the length of a rectangle is 50ft. The perimeter of the rectangle is 275ft more than the width of the rectangle. find the width. I'm stumped, it's been years since i tried this and I feel stupid lol thanks for the help
Answer by ankor@dixie-net.com(22740)

the length of a rectangle is 50ft.
The perimeter of the rectangle is 275ft more than the width of the rectangle.
find the width.
:
Let W = the width
:
The perimeter equation
2L + 2W = 275 + W
we know the length is 50, so twice that, we have
100 + 2W = 275 + W
2W - W = 275 - 100
W = 175 ft is the width
:
:
Check that, the perimeter = 175 + 275 = 450 ft
2(50) + 2(175) =
100 + 350 = 450